OpenStudy (veeveeniko):

How would I write the equation of a circle with a center at (0, 0) and a radius of 5?

OpenStudy (anonymous):

x^2+y^2=r^2 x^2+y^2=5^2 x^2+y^2=25 http://www.regentsprep.org/Regents/math/algtrig/ATC1/circlelesson.htm